# gPC-based robustness analysis of neural systems through probabilistic recurrence metrics

### v1

@@ 2.2 Generalised polynomial chaos (gPC) surrogate models
To compute statistics of the stochastic process $\left\{\mathbf{x}\left(t;Z\right)\right\}_{t\geq 0}$ described by (1), we employ a gPC approximation [xiu2003modeling] to obtain a surrogate model of (1), based on polynomial regression. The gPC surrogate model is simpler to integrate and retains the probabilistic information needed to efficiently calculate the mean process $\left\{\mathbb{E}\left[\mathbf{x}\left(t;Z\right)\right]=:\mathbb{E}[\mathbf{x}](t)\right\}_{t\geq 0}$, where $\mathbb{E}$ denotes the expectation with respect to $Z\sim\mathcal{U}(\mathcal{A})$. This gPC-based approach has been recently shown to provide surrogate models characterised by a favourable trade-off between computational effort and accuracy, for relevant systems in neuroscience at different spatial scales [sutulovic2024efficient].
@@
We assume that the stochastic process $\left\{\mathbf{x}(t;Z)\right\}_{t\geq 0}$ has finite variance, which is not restrictive for mathematical models in neuroscience, since any meaningful model would take into account biophysical constraints, thus producing bounded state trajectories. Specifically, we assume that for each $t\geq 0$ (outside perhaps of a zero measure set), every component of $\mathbf{x}(t;\cdot)$, treated as a function of $Z$, belongs to the space $L_{\mathcal{U}}^{2}(\mathcal{A})=\{\psi\colon\mathbb{E}[\psi^{2}]:=\frac{1}{\text{size}(\mathcal{A})}\int_{A}\psi^{2}(\xi)d\xi<+\infty\}$. This is guaranteed if the function $\mathbf{x}(t;\cdot)$ is measurable and bounded. Since the components of $Z$ are independent, a natural basis for $L_{\mathcal{U}}^{2}(\mathcal{A})$ is the set of multivariable polynomials $\{\Phi_{\alpha}\}_{\alpha}$ that satisfy the orthogonality relation $\mathbb{E}[\Phi_{\alpha}\Phi_{\beta}]=\frac{1}{\text{size}(\mathcal{A})}\int_{A}\Phi_{\alpha}(\xi)\Phi_{\beta}(\xi)d\xi=\delta_{\alpha\beta}$, where $\delta_{\alpha\beta}$ is the $d$-variate Kronecker delta and $\alpha=(\alpha_{1},\ldots,\alpha_{d})\in\mathbb{N}_{0}^{d}$ is a multi-index. The basis of orthogonal polynomials $\{\Phi_{\alpha}\}_{\alpha}$ can be computed via a Gram-Schmidt orthogonalisation process [xiu2003modeling]. As proven in [cameron1947orthogonal], the following series representation holds in $L^{2}_{\mathcal{U}}(\mathcal{A})$:
@@
where the choice of the basis $\{\Phi_{\alpha}\}_{\alpha}$ depends on the probability distribution of the random variables in $Z$, according to the Wiener-Askey table [xiu2002wiener, xiu2003modeling] that associates probability distributions with classes of orthogonal polynomials. In our case, since the random variables in $Z$ are uniformly distributed, $\{\Phi_{\alpha}\}_{\alpha}$ is composed of Legendre polynomials. Importantly, the series representation (4) neatly decouples the effect of the uncertain parameters $Z$, confined to the basis of orthogonal polynomials $\{\Phi_{\alpha}(Z)\}_{\alpha}$, and the deterministic time dependence, contained in the spectral coefficients $\{\mathbf{x}_{\alpha}(t)\}_{\alpha}$. This greatly simplifies the computation of statistics of the random process $\mathbf{x}(t;Z)$; for instance, mean and variance are given by
@@
To obtain a tractable surrogate model for practical implementation (Figure 1b), we truncate the series (4) to an expansion order $M$, namely, we only retain indices with $|\alpha|:=\sum_{i=1}^{d}\alpha_{i}\leq M$, thus approximating $\mathbf{x}(t;Z)$ with $\mathbf{x}_{M}(t;Z)=\sum_{|\alpha|\leq M}\mathbf{x}_{\alpha}(t)\Phi_{\alpha}(Z)$. Therefore, construction of the surrogate model only requires finding the spectral coefficients $\{\mathbf{x}_{\alpha}\}_{|\alpha|\leq M}$. For this, we adopt the non-intrusive collocation approach, which consists of computing $S_{\text{C}}\geq\binom{M+d}{d}$ system trajectory realisations and then estimating the spectral coefficients indirectly, e.g., via a least-squares approximation [blatman2011adaptive, sutulovic2024efficient]. Since theoretical convergence guarantees are available for static maps of random variables [xiu2010numerical], but not for solutions to general ODE systems, it is important to numerically check the suitability of the gPC-based approximation, by computing the coefficients $\{\mathbf{x}_{\alpha}(t)\}_{\left|\alpha\right|\leq M}$ with increasing values of $M$, until the decimal expansion of a fixed initial set of coefficients $\{\mathbf{x}_{\alpha}(t)\}_{\left|\alpha\right|\leq M_{0}}$, with $M_{0}<M$, stabilizes to a prescribed order of decimal digits. For computational purposes, we implement gPC collocation via the PoCET toolbox in Matlab [Petzke2020_PoCET].
@@ 2.3 Mean state trajectories and output signals
Sampling the mean system output yields a time-series $\{\mathbb{E}[y](t_{k})\}_{k}$, formed by a collection of average values at time instants $\{t_{k}\}_{k}=\{k\cdot\Delta t\}_{k}$, $k\in\mathbb{N}$, where $\Delta t>0$ is a fixed inter-sampling time. Then, we need to design criteria to determine whether, in the presence of the uncertainty $Z$, the sampled mean output signal $\{\mathbb{E}[y](t_{k})\}_{k}$ preserves the qualitative features that characterise the nominal regime. Assessing regime preservation in the presence of probabilistic parametric uncertainty is particularly challenging: we cannot resort to the standard metrics employed in the deterministic case for regime identification, as for mean signals they either cannot be computed or lose their interpretability (see Supplementary Section S2). In the deterministic case, i.e., for fixed parameter values, standard simple metrics such as inter-spike interval, duty cycle or spike count, often employed to extract quantitative information from neural signals [barrio2011parameter, izhikevich2007dynamical, jirsa2014nature], are easy to compute. However, with probabilistic uncertainty, the mean output signal is obtained by averaging different realisations of the output signal, whose spikes and bursts occur at different time instants or intervals. Hence, standard metrics are not properly defined for mean signals, where it is not even clear what constitutes a spike or a burst (see Supplementary Figure S5). Thresholding procedures further hamper the generic use of such standard metrics [pavlov2000extracting]. To distinguish between different regimes, an alternative approach may be to consider the mean output signal as a time-series and compute its power spectral density (or some wavelet transform) to search for distinctive frequency components [rosso2006eeg, meisel2012scaling] (Supplementary Figure S9). However, through numerical investigations partially reported in Supplementary Section S2, we verified the current inefficacy of these and other existing methods used for regime identification in deterministic models, which become generally inapplicable in the presence of probabilistic parametric uncertainties.
In deterministic realisations of the HR model, the plateau-bursting regime (D) in Figure 2 is characterised by variable $x_{1}$ exhibiting oscillations, initiated after it reaches the value $x_{1}\approx 0$, which last with gradually damped amplitude for a finite time interval, after which $x_{1}\approx 0$ again. If such a distinctive characteristic is also present in the mean output signal for a given parametric uncertainty (as in Figure 1c), then we may say that the regime is preserved in a probabilistic sense [sutulovic2024efficient, Section III.C]. Unfortunately, cases in which the mean signal preserves such properties of the corresponding nominal regime are extremely rare. For the HR model, this happens only for plateau-bursting and not for other regimes; for other systems, it may never happen.
@@ 2.4.1 Recurrence analysis
To extract information from mean output signals aimed at assessing regime preservation under parametric uncertainty, we resort to recurrence analysis [lopes2021recurrence]. An advantage of recurrence plots [eckmann1995recurrence] lies in their ability to effectively visualise characteristic features of signals through constructed geometric patterns, which reveal qualitative signatures of, e.g., periodic and quasi-periodic behaviour, slowly evolving dynamics, stochastic-like fluctuations, or deterministic chaotic regimes [marwan2007recurrence, prado2020maximum]. To construct the recurrence plot, given $T$ samples of the mean output $\{\mathbb{E}[y](t_{k})\}_{k=1}^{T}$ as our discrete-time signal of interest, we compute the distance matrix $D$ as
@@
The entries of $D$ in (6) yield an unthresholded recurrence plot [marwan2007recurrence, Subsection 3.2.5], which processes the distance matrix $D$ as an image with pixels coloured according to the values of its entries. Since the recurrence patterns are preserved under scaling, without loss of generality, $D$ is normalised by its maximum value, so that all of its entries $D_{\ell h}$ lie in $[0,1]$. The recurrence plots for the HR model (2), corresponding to distinct deterministic realisations of $y(t)=x_{1}(t)$, the membrane potential, in different dynamical regimes, are shown in Figure 2 (bottom); see also Figure 1d, left. Recurrence plots allow us to visually distinguish between different dynamical regimes in a clear manner. This is also true for the JR model (3), see Supplementary Figure S2.
@@
Recurrence analysis allow us to transform the problem of assessing regime preservation into image pattern recognition/retention. This task is often tackled using neural networks [ripley2007pattern], which however are not suitable in our case, because training and validation processes do not generalise well from one system to another, require extensive labelling, and are computationally heavy. A possible alternative would be unsupervised classification of images [gadetsky2024let], which however requires knowing the number of labels in advance; since we are not tackling the problem of global classification of dynamical regimes in a parameter space, and we are only interested in a local classification aimed at assessing local regime preservation under parametric uncertainty, we take a more direct and interpretable approach that is not based on machine learning algorithms, and leave their investigation for future research.
@@ 2.4.2 Automated blob counting for recurrence plots
On a side note, the most suitable threshold $\vartheta$ to be used is not known a priori and likely depends on the specific dynamical regime. As described in Figure 4, we drew inspiration for threshold selection from persistent homology [chazal2021introduction], which tracks how topological structures (such as connected components or holes) form and disappear, and employs such changes to distinguish between enduring patterns, as opposed to short-lived noisy structures, in data. Similarly, we study how the blob count varies or persists through all possible thresholds $\vartheta\in(0,1)$. The most persistent blob count can be different from the most frequent blob count; thus, the employed persistence bar charts differ from simple histograms.
@@ 3.1 Robustness of bursting at the single-neuron level
We construct PRP plots for the single-neuron HR model (2) from [hindmarsh1984model] and focus on the preservation of bursting activity in the membrane potential $x_{1}(t)$ under parametric uncertainty. Our interest is driven by the fact that bursting is a fundamental element of neural computations, to reliably transmit messages to post-synaptic neurons and overcome transmission failures [lisman1997bursts], and enable selective communication at the inter-neuron level [izhikevich2003bursts]. Bursting is also important in triggering insulin secretion in $\beta$ cells within a pancreatic islet [beauvois2006glucose] and release of growth hormone in cultured pituitary somatrophs [tsaneva2007mechanism], in sensory processing in the thalamocortical relay [sherman2002role], and in ensuring temporal summation and long-term potentiation [lisman1997bursts]. So, understanding how much parametric uncertainty the system can withstand while preserving bursting is relevant for many vital biological functions. To this end, we apply our methodology, described in Section 2, to $\mathbb{E}[x_{1}](t)$. Studying mean membrane potentials is justified by the principle of averaging [van1992theory, faisal2008noise], a natural mechanism in which the same noisy stimulus is processed in parallel by a redundant group of neurons, whose outputs are then averaged in the post-synaptic neurons, leading to a more fault-tolerant computation. Examples of averaging include neighbouring retinal ganglion cells processing visual signals [puchalla2005redundancy] and stereocilia of auditory cells processing sounds [kozlov2007coherent].
@@
Figure 6: PRA of the bursting activity of membrane potential $x_{1}$ in the HR model (2) under probabilistic parametric uncertainty, in the parameter spaces $\mathcal{P}_{1}=[b_{\min},b_{\max}]\times[I_{\min},I_{\max}]=[2.5,3.3]\times[2.2,4.4]$ with $r=0.01$ (top) and $\mathcal{P}_{2}=[r_{\min},r_{\max}]\times[I_{\min},I_{\max}]=[0.005,0.03]\times[2.4,3.4]$ with $b=3$ (bottom). Mean signals are computed through gPC surrogate models via the collocation method, with maximum expansion order $M=5$ and $S_{\text{C}}=250$ collocation samples. In $\mathcal{P}_{1}$, $\Delta b_{\max}=0.1$ and $\Delta I_{\max}=0.2$; in $\mathcal{P}_{2}$, $\Delta r_{\max}=0.005$ and $\Delta I_{\max}=0.1$. In both cases, $N=8$ uncertainty levels are considered. In $\mathcal{P}_{1}$, we show: (a1) deterministic dynamical regime classification based on bifurcation analysis, using spike count as a metric (image from [barrio2011parameter], courtesy of Springer); (b1) PRP plot obtained with the LE approach; (c1) PRP plot obtained with the FG approach, on a grid of $27\times 36$ equidistant points. In $\mathcal{P}_{2}$, we show: (a2) deterministic dynamical regime classification based on the largest Lyapunov exponent, which only highlights regions associated with the chaotic regime (image from [gu2013biological], courtesy of PLoS One); (b2) PRP plot obtained with the LE approach; (c2) PRP plot obtained with the FG approach, on a grid of $30\times 18$ equidistant points.
@@
We show the PRP plots computed for $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ with the LE approach in Figure 6b1 and b2, respectively, and with the FG approach in Figure 6c1 and c2 respectively. The plots in Figure 6a1 [barrio2011parameter] and a2 [gu2013biological] display, in the same parameter spaces as the PRP plots, the regimes resulting from different fixed parameter values in a purely deterministic setting, and hence offer an intrinsically different and complementary perspective. The PRP plots visualise the preservation of dynamical regimes, identified through the blob count for the recurrence plots of the mean membrane potential $\mathbb{E}[x_{1}]$, in the probabilistic setting of (1). The blob count is encoded either in the colour of the rectangles in Figure 6b1 and b2, or in the colour map in the left plots in Figure 6c1 and c2. The probabilistic robustness of nominal regimes in the presence of uniformly distributed parametric uncertainty is quantified either through the size of the rectangles in Figure 6b1 and b2, or through the preservation percentage in the right plots in Figure 6c1 and c2.
@@
In the parameter space $\mathcal{P}_{1}$, the results stemming from our methodology reveal a “gradient” of colours associated with increasing blob counts, which can be observed both in Figure 6b1 (LE approach) and Figure 6c1, left (FG approach). Although we do not know a priori which regime corresponds to each colour, comparison with the bifurcation plot in Figure 6a1 [barrio2011parameter], which conveys information on the deterministic nominal regime associated with each point in the parameter space, allows us to identify the regime corresponding to a certain colour in the PRP plot: in Figure 6b1 and Figure 6c1, left, we can associate the upper-left light blue region with plateau bursting, the shades from turquoise to olive green to orange with square-wave bursting, and the yellow region with spiking, while the dark blue area encompasses both the chaotic and the quiescence regions, as discussed in depth later in this section. The probabilistic robustness of each regime is quantified by the size of the rectangles in Figure 6b1 (the longer the side of the rectangle, the more robust the system is with respect to the corresponding parameter) and by the preservation percentage plot in Figure 6c1, right (the darker the colour, the higher the preservation percentage). In both cases, plateau bursting exhibits more robustness to parametric uncertainty than square-wave bursting, even though it is nominally emerging in a smaller portion of the parameter space $\mathcal{P}_{1}$. These findings are consistent with the robustness properties of bursting observed in a Hodgkin-Huxley-like model of pancreatic islets in [smolen1993pancreatic].
@@
In the parameter space $\mathcal{P}_{2}$, the deterministic plot based on Lyapunov exponents in Figure 6a2 [gu2013biological] highlights the coloured regions associated with a chaotic regime, which emerge from the blue background. In our PRP plots, these regions have zero blob count (dark blue in Figure 6b2 and in Figure 6c2, left) and are particularly robust to parametric uncertainty (in Figure 6b2, many dark blue rectangles have maximal area given by the user-defined choice; in Figure 6c2, right, the regions have high preservation percentage). However, a zero blob count also characterises the quiescence regime observed in the parameter space $\mathcal{P}_{1}$: also the quiescent dark blue areas in Figure 6b1 and in Figure 6c1, left, are very robust (large rectangles in Figure 6b1 and high preservation percentage in Figure 6c1, right). In fact, discerning meaningful patterns from mean chaotic state trajectories in the presence of parametric uncertainty is particularly challenging: in a chaotic regime, the mean membrane potential computed in a probabilistic setting is indistinguishable from a zero signal affected by random noise, because it results from averaging deterministic realisations of firing patterns of the membrane potential that are very dissimilar, even for parameter values that are numerically close. This yields a rapid degradation of the recurrence patterns and hence a zero blob count, even for low uncertainty levels, thus making the mean signal in the chaotic regime indistinguishable from that in the quiescence regime, as is apparent from Figure 6b1, where no discontinuity is visible between the chaotic bursting and the quiescence regions (for another example, see Supplementary Figure S4). On the other hand, chaos disrupts the probabilistic robustness of the bursting regimes even for small uncertainty levels, as shown by the fragmented tiling on the diagonal of Figure 6b1, on the border between square-wave bursting and chaotic bursting regions (see also Supplementary Figure S16). Our PRP plots also reveal other, non-chaotic regions in $\mathcal{P}_{2}$, corresponding to large blob counts, where however the associated regime has a low level of robust preservation under uncertainty.

### v2

@@ 2.2 Generalised polynomial chaos (gPC) surrogate models
To compute statistics of the stochastic process $\left\{\mathbf{x}\left(t;Z\right)\right\}_{t\geq 0}$ described by (1), we employ a gPC approximation [74] to obtain a surrogate model of (1), based on polynomial regression. The gPC surrogate model is simpler to integrate and retains the probabilistic information needed to efficiently calculate the mean process $\left\{\mathbb{E}\left[\mathbf{x}\left(t;Z\right)\right]=:\mathbb{E}[\mathbf{x}](t)\right\}_{t\geq 0}$, where $\mathbb{E}$ denotes the expectation with respect to $Z\sim\mathcal{U}(\mathcal{A})$. This gPC-based approach has been recently shown to provide surrogate models characterised by a favourable trade-off between computational effort and accuracy, for relevant systems in neuroscience at different spatial scales [69].
@@
We assume that the stochastic process $\left\{\mathbf{x}(t;Z)\right\}_{t\geq 0}$ has finite variance, which is not restrictive for mathematical models in neuroscience, since any meaningful model would take into account biophysical constraints, thus producing bounded state trajectories. Specifically, we assume that for each $t\geq 0$ (outside perhaps of a zero measure set), every component of $\mathbf{x}(t;\cdot)$, treated as a function of $Z$, belongs to the space $L_{\mathcal{U}}^{2}(\mathcal{A})=\{\psi\colon\mathbb{E}[\psi^{2}]:=\frac{1}{\text{size}(\mathcal{A})}\int_{A}\psi^{2}(\xi)d\xi<+\infty\}$. This is guaranteed if the function $\mathbf{x}(t;\cdot)$ is measurable and bounded. Since the components of $Z$ are independent, a natural basis for $L_{\mathcal{U}}^{2}(\mathcal{A})$ is the set of multivariable polynomials $\{\Phi_{\alpha}\}_{\alpha}$ that satisfy the orthogonality relation $\mathbb{E}[\Phi_{\alpha}\Phi_{\beta}]=\frac{1}{\text{size}(\mathcal{A})}\int_{A}\Phi_{\alpha}(\xi)\Phi_{\beta}(\xi)d\xi=\delta_{\alpha\beta}$, where $\delta_{\alpha\beta}$ is the $d$-variate Kronecker delta and $\alpha=(\alpha_{1},\ldots,\alpha_{d})\in\mathbb{N}_{0}^{d}$ is a multi-index. The basis of orthogonal polynomials $\{\Phi_{\alpha}\}_{\alpha}$ can be computed via a Gram-Schmidt orthogonalisation process [74]. As proven in [13], the following series representation holds in $L^{2}_{\mathcal{U}}(\mathcal{A})$:
@@
where the choice of the basis $\{\Phi_{\alpha}\}_{\alpha}$ depends on the probability distribution of the random variables in $Z$, according to the Wiener-Askey table [73, 74] that associates probability distributions with classes of orthogonal polynomials. In our case, since the random variables in $Z$ are uniformly distributed, $\{\Phi_{\alpha}\}_{\alpha}$ is composed of Legendre polynomials. Importantly, the series representation (4) neatly decouples the effect of the uncertain parameters $Z$, confined to the basis of orthogonal polynomials $\{\Phi_{\alpha}(Z)\}_{\alpha}$, and the deterministic time dependence, contained in the spectral coefficients $\{\mathbf{x}_{\alpha}(t)\}_{\alpha}$. This greatly simplifies the computation of statistics of the random process $\mathbf{x}(t;Z)$; for instance, mean and variance are given by
@@
To obtain a tractable surrogate model for practical implementation (Figure 1b), we truncate the series (4) to an expansion order $M$, namely, we only retain indices with $|\alpha|:=\sum_{i=1}^{d}\alpha_{i}\leq M$, thus approximating $\mathbf{x}(t;Z)$ with $\mathbf{x}_{M}(t;Z)=\sum_{|\alpha|\leq M}\mathbf{x}_{\alpha}(t)\Phi_{\alpha}(Z)$. Therefore, construction of the surrogate model only requires finding the spectral coefficients $\{\mathbf{x}_{\alpha}\}_{|\alpha|\leq M}$. For this, we adopt the non-intrusive collocation approach, which consists of computing $S_{\text{C}}\geq\binom{M+d}{d}$ system trajectory realisations and then estimating the spectral coefficients indirectly, e.g., via a least-squares approximation [11, 69]. Since theoretical convergence guarantees are available for static maps of random variables [75], but not for solutions to general ODE systems, it is important to numerically check the suitability of the gPC-based approximation, by computing the coefficients $\{\mathbf{x}_{\alpha}(t)\}_{\left|\alpha\right|\leq M}$ with increasing values of $M$, until the decimal expansion of a fixed initial set of coefficients $\{\mathbf{x}_{\alpha}(t)\}_{\left|\alpha\right|\leq M_{0}}$, with $M_{0}<M$, stabilizes to a prescribed order of decimal digits. For computational purposes, we implement gPC collocation via the PoCET toolbox in Matlab [56].
@@ 2.3 Mean state trajectories and output signals
Sampling the mean system output yields a time-series $\{\mathbb{E}[y](t_{k})\}_{k}$, formed by a collection of average values at time instants $\{t_{k}\}_{k}=\{k\cdot\Delta t\}_{k}$, $k\in\mathbb{N}$, where $\Delta t>0$ is a fixed inter-sampling time. Then, we need to design criteria to determine whether, in the presence of the uncertainty $Z$, the sampled mean output signal $\{\mathbb{E}[y](t_{k})\}_{k}$ preserves the qualitative features that characterise the nominal regime. Assessing regime preservation in the presence of probabilistic parametric uncertainty is particularly challenging: we cannot resort to the standard metrics employed in the deterministic case for regime identification, as for mean signals they either cannot be computed or lose their interpretability (see Supplementary Section S2). In the deterministic case, i.e., for fixed parameter values, standard simple metrics such as inter-spike interval, duty cycle or spike count, often employed to extract quantitative information from neural signals [7, 29, 31], are easy to compute. However, with probabilistic uncertainty, the mean output signal is obtained by averaging different realisations of the output signal, whose spikes and bursts occur at different time instants or intervals. Hence, standard metrics are not properly defined for mean signals, where it is not even clear what constitutes a spike or a burst (see Supplementary Figure S5). Thresholding procedures further hamper the generic use of such standard metrics [54]. To distinguish between different regimes, an alternative approach may be to consider the mean output signal as a time-series and compute its power spectral density (or some wavelet transform) to search for distinctive frequency components [62, 47] (Supplementary Figure S9). However, through numerical investigations partially reported in Supplementary Section S2, we verified the current inefficacy of these and other existing methods used for regime identification in deterministic models, which become generally inapplicable in the presence of probabilistic parametric uncertainties.
In deterministic realisations of the HR model, the plateau-bursting regime (D) in Figure 2 is characterised by variable $x_{1}$ exhibiting oscillations, initiated after it reaches the value $x_{1}\approx 0$, which last with gradually damped amplitude for a finite time interval, after which $x_{1}\approx 0$ again. If such a distinctive characteristic is also present in the mean output signal for a given parametric uncertainty (as in Figure 1c), then we may say that the regime is preserved in a probabilistic sense [69, Section III.C]. Unfortunately, cases in which the mean signal preserves such properties of the corresponding nominal regime are extremely rare. For the HR model, this happens only for plateau-bursting and not for other regimes; for other systems, it may never happen.
@@ 2.4.1 Recurrence analysis
To extract information from mean output signals aimed at assessing regime preservation under parametric uncertainty, we resort to recurrence analysis [42]. An advantage of recurrence plots [15] lies in their ability to effectively visualise characteristic features of signals through constructed geometric patterns, which reveal qualitative signatures of, e.g., periodic and quasi-periodic behaviour, slowly evolving dynamics, stochastic-like fluctuations, or deterministic chaotic regimes [46, 58]. To construct the recurrence plot, given $T$ samples of the mean output $\{\mathbb{E}[y](t_{k})\}_{k=1}^{T}$ as our discrete-time signal of interest, we compute the distance matrix $D$ as
@@
The entries of $D$ in (6) yield an unthresholded recurrence plot [46, Subsection 3.2.5], which processes the distance matrix $D$ as an image with pixels coloured according to the values of its entries. Since the recurrence patterns are preserved under scaling, without loss of generality, $D$ is normalised by its maximum value, so that all of its entries $D_{\ell h}$ lie in $[0,1]$. The recurrence plots for the HR model (2), corresponding to distinct deterministic realisations of $y(t)=x_{1}(t)$, the membrane potential, in different dynamical regimes, are shown in Figure 2 (bottom); see also Figure 1d, left. Recurrence plots allow us to visually distinguish between different dynamical regimes in a clear manner. This is also true for the JR model (3), see Supplementary Figure S2.
@@
Recurrence analysis allow us to transform the problem of assessing regime preservation into image pattern recognition/retention. This task is often tackled using neural networks [61], which however are not suitable in our case, because training and validation processes do not generalise well from one system to another, require extensive labelling, and are computationally heavy. A possible alternative would be unsupervised classification of images [18], which however requires knowing the number of labels in advance; since we are not tackling the problem of global classification of dynamical regimes in a parameter space, and we are only interested in a local classification aimed at assessing local regime preservation under parametric uncertainty, we take a more direct and interpretable approach that is not based on machine learning algorithms, and leave their investigation for future research.
@@ 2.4.2 Automated blob counting for recurrence plots
On a side note, the most suitable threshold $\vartheta$ to be used is not known a priori and likely depends on the specific dynamical regime. As described in Figure 4, we drew inspiration for threshold selection from persistent homology [14], which tracks how topological structures (such as connected components or holes) form and disappear, and employs such changes to distinguish between enduring patterns, as opposed to short-lived noisy structures, in data. Similarly, we study how the blob count varies or persists through all possible thresholds $\vartheta\in(0,1)$. The most persistent blob count can be different from the most frequent blob count; thus, the employed persistence bar charts differ from simple histograms.
@@ 3.1 Robustness of bursting at the single-neuron level
We construct PRP plots for the single-neuron HR model (2) from [26] and focus on the preservation of bursting activity in the membrane potential $x_{1}(t)$ under parametric uncertainty. Our interest is driven by the fact that bursting is a fundamental element of neural computations, to reliably transmit messages to post-synaptic neurons and overcome transmission failures [41], and enable selective communication at the inter-neuron level [28]. Bursting is also important in triggering insulin secretion in $\beta$ cells within a pancreatic islet [8] and release of growth hormone in cultured pituitary somatrophs [71], in sensory processing in the thalamocortical relay [63], and in ensuring temporal summation and long-term potentiation [41]. So, understanding how much parametric uncertainty the system can withstand while preserving bursting is relevant for many vital biological functions. To this end, we apply our methodology, described in Section 2, to $\mathbb{E}[x_{1}](t)$. Studying mean membrane potentials is justified by the principle of averaging [72, 16], a natural mechanism in which the same noisy stimulus is processed in parallel by a redundant group of neurons, whose outputs are then averaged in the post-synaptic neurons, leading to a more fault-tolerant computation. Examples of averaging include neighbouring retinal ganglion cells processing visual signals [60] and stereocilia of auditory cells processing sounds [37].
@@
Figure 6: PRA of the bursting activity of membrane potential $x_{1}$ in the HR model (2) under probabilistic parametric uncertainty, in the parameter spaces $\mathcal{P}_{1}=[b_{\min},b_{\max}]\times[I_{\min},I_{\max}]=[2.5,3.3]\times[2.2,4.4]$ with $r=0.01$ (top) and $\mathcal{P}_{2}=[r_{\min},r_{\max}]\times[I_{\min},I_{\max}]=[0.005,0.03]\times[2.4,3.4]$ with $b=3$ (bottom). Mean signals are computed through gPC surrogate models via the collocation method, with maximum expansion order $M=5$ and $S_{\text{C}}=250$ collocation samples. In $\mathcal{P}_{1}$, $\Delta b_{\max}=0.1$ and $\Delta I_{\max}=0.2$; in $\mathcal{P}_{2}$, $\Delta r_{\max}=0.005$ and $\Delta I_{\max}=0.1$. In both cases, $N=8$ uncertainty levels are considered. In $\mathcal{P}_{1}$, we show: (a1) deterministic dynamical regime classification based on bifurcation analysis, using spike count as a metric (image from [7], courtesy of Springer); (b1) PRP plot obtained with the LE approach; (c1) PRP plot obtained with the FG approach, on a grid of $27\times 36$ equidistant points. In $\mathcal{P}_{2}$, we show: (a2) deterministic dynamical regime classification based on the largest Lyapunov exponent, which only highlights regions associated with the chaotic regime (image from [23], courtesy of PLoS One); (b2) PRP plot obtained with the LE approach; (c2) PRP plot obtained with the FG approach, on a grid of $30\times 18$ equidistant points.
@@
We show the PRP plots computed for $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ with the LE approach in Figure 6b1 and b2, respectively, and with the FG approach in Figure 6c1 and c2 respectively. The plots in Figure 6a1 [7] and a2 [23] display, in the same parameter spaces as the PRP plots, the regimes resulting from different fixed parameter values in a purely deterministic setting, and hence offer an intrinsically different and complementary perspective. The PRP plots visualise the preservation of dynamical regimes, identified through the blob count for the recurrence plots of the mean membrane potential $\mathbb{E}[x_{1}]$, in the probabilistic setting of (1). The blob count is encoded either in the colour of the rectangles in Figure 6b1 and b2, or in the colour map in the left plots in Figure 6c1 and c2. The probabilistic robustness of nominal regimes in the presence of uniformly distributed parametric uncertainty is quantified either through the size of the rectangles in Figure 6b1 and b2, or through the preservation percentage in the right plots in Figure 6c1 and c2.
@@
In the parameter space $\mathcal{P}_{1}$, the results stemming from our methodology reveal a “gradient” of colours associated with increasing blob counts, which can be observed both in Figure 6b1 (LE approach) and Figure 6c1, left (FG approach). Although we do not know a priori which regime corresponds to each colour, comparison with the bifurcation plot in Figure 6a1 [7], which conveys information on the deterministic nominal regime associated with each point in the parameter space, allows us to identify the regime corresponding to a certain colour in the PRP plot: in Figure 6b1 and Figure 6c1, left, we can associate the upper-left light blue region with plateau bursting, the shades from turquoise to olive green to orange with square-wave bursting, and the yellow region with spiking, while the dark blue area encompasses both the chaotic and the quiescence regions, as discussed in depth later in this section. The probabilistic robustness of each regime is quantified by the size of the rectangles in Figure 6b1 (the longer the side of the rectangle, the more robust the system is with respect to the corresponding parameter) and by the preservation percentage plot in Figure 6c1, right (the darker the colour, the higher the preservation percentage). In both cases, plateau bursting exhibits more robustness to parametric uncertainty than square-wave bursting, even though it is nominally emerging in a smaller portion of the parameter space $\mathcal{P}_{1}$. These findings are consistent with the robustness properties of bursting observed in a Hodgkin-Huxley-like model of pancreatic islets in [65].
@@
In the parameter space $\mathcal{P}_{2}$, the deterministic plot based on Lyapunov exponents in Figure 6a2 [23] highlights the coloured regions associated with a chaotic regime, which emerge from the blue background. In our PRP plots, these regions have zero blob count (dark blue in Figure 6b2 and in Figure 6c2, left) and are particularly robust to parametric uncertainty (in Figure 6b2, many dark blue rectangles have maximal area given by the user-defined choice; in Figure 6c2, right, the regions have high preservation percentage). However, a zero blob count also characterises the quiescence regime observed in the parameter space $\mathcal{P}_{1}$: also the quiescent dark blue areas in Figure 6b1 and in Figure 6c1, left, are very robust (large rectangles in Figure 6b1 and high preservation percentage in Figure 6c1, right). In fact, discerning meaningful patterns from mean chaotic state trajectories in the presence of parametric uncertainty is particularly challenging: in a chaotic regime, the mean membrane potential computed in a probabilistic setting is indistinguishable from a zero signal affected by random noise, because it results from averaging deterministic realisations of firing patterns of the membrane potential that are very dissimilar, even for parameter values that are numerically close. This yields a rapid degradation of the recurrence patterns and hence a zero blob count, even for low uncertainty levels, thus making the mean signal in the chaotic regime indistinguishable from that in the quiescence regime, as is apparent from Figure 6b1, where no discontinuity is visible between the chaotic bursting and the quiescence regions (for another example, see Supplementary Figure S4). On the other hand, chaos disrupts the probabilistic robustness of the bursting regimes even for small uncertainty levels, as shown by the fragmented tiling on the diagonal of Figure 6b1, on the border between square-wave bursting and chaotic bursting regions (see also Supplementary Figure S16). Our PRP plots also reveal other, non-chaotic regions in $\mathcal{P}_{2}$, corresponding to large blob counts, where however the associated regime has a low level of robust preservation under uncertainty.